I do not understand how to solve this problem and am not grasping the problem solving steps. I am not sure where to start or what to add or subtract or anything. It is like a foreign language! I need someone to explain what the steps are and where to start.
Objectives are Solve two step equations (integers) and Solve two step equations (decimal and fraction)
-9x+11/24=-19/24

Found 2 solutions by rfer, edjones:
Answer by rfer(16322)

-9x+11/24=-19/24
subtract 11/24
-9x=-19/24-11/24
-9x=-30/24
divide by -9
x=(-30/24)/-9
invert and multiply
x=-30/24*-1/9
x=+30/216
x=5/36

Answer by edjones(8007)

-9x + 11/24 = -19/24
-216x+11=-19 Multiply each side by 24.
-216x=-30
x=5/36
.
Ed